# PAS: Data-Efficient Plug-and-Play Prompt Augmentation System

### v1

@@ 3.1. Prompts Data Collection
In this section, we introduce the prompt collection process. To generate high-quality (prompt, complementary prompt) pairs, we first need to select high-quality prompts. To achieve this, we use two curated datasets: the LMSYS-1M dataset (Zheng et al., 2023) and the WildChat dataset (Zhao et al., 2024). As shown in Figure 3(a), our data selection process involves three main steps to ensure the quality and relevance of the data. First, we perform deduplication using the SimCSE bge model to embed the prompts, followed by the HNSW clustering algorithm to group these embeddings. From each cluster, we extract a small amount of data to reduce redundancy. Second, we conduct quality filtering using the BaiChuan 13b model to score the data and filter out low-quality entries, thereby enhancing the overall data quality. Finally, we perform classification. We utilize 60,000 internally labeled classification data points from BaiChuan to fine-tune (SFT) a BaiChuan 13b (Yang et al., 2023b) model, resulting in a classification model that we then use to categorize the prompts into commonly used categories such as Q&A and coding. This series of steps ensures the diversity, quality, and accurate classification of the data.
@@ 21 $isCorrect\leftarrow\text{FewShot Eval with }D_{\text{golden}}$;
To generate a high-quality prompt complementary dataset, we designed an automated data generation pipeline based on Few-Shot Learning. The algorithm mainly consists of two phases: data generation and data selection with regeneration. We use a set of golden data $D_{\text{golden}}=\{(p_{i},a_{i})\}_{i=1}^{N}$, containing 4 to 5 pairs of few-shot examples for each category from BaiChuan. Then, we utilize the prompt dataset $P$ from Section 3.1 to generate high-quality (prompt, complementary prompt) pairs.
@@
For each prompt $p_{j}\in P$ in every category, we utilize the Few-Shot Learning method based on the prompt in Figure 4 to generate a corresponding complementary prompt $a_{j}$ based on the golden data $D_{\text{golden}}$. The generated prompt-complementary prompt pair $(p_{j},a_{j})$ is then added to $D_{\text{generated}}$.
@@ 3.3. Prompt Complementary Dataset
In this section, we provide a detailed analysis of the generated prompt-complementary dataset. The prompt-complementary dataset consists of approximately 9,000 high-quality (prompt, complementary prompt) pairs. We summarize the dataset distribution in Figure 7. From Figure 7, it is evident that the data are categorized into 14 categories, with each category containing about 500 data points. This broad coverage across various prompt cases demonstrates the dataset’s exceptional generalization ability. Additionally, there is a substantial amount of Coding and Q&A data. Given that Coding and Q&A are among the most commonly used functions, this distribution is reasonable and emphasizes the most critical functionalities.
Despite these advantages in data distribution, our automatic complementary prompt data generation process described in Section 3.2 allows us to control the categories of the generated data. This flexibility enables our method to cater to both general-purpose models and to generate specialized data to enhance prompt capabilities in specific domains.
@@ 4.1. Experiments Setting
For PAS models, we select several smaller models to efficiently train a PAS model. We select Qwen2-7b-Chat (Bai et al., 2023), LLaMA-2-7b-instruct (Touvron et al., 2023). Then we utilize our trained PAS models to some massive SoTA models, i.e. GPT-4-turbo-2024-04-09, GPT-4-1106-preview, GPT-4-0613, GPT-3.5-turbo-1106 (OpenAI, 2023a), Qwen2-72B-Chat (Bai et al., 2023), and LLaMA-3-70b-instruct (Touvron et al., 2023; meta llama, 2024).
@@
For Qwen2-7B-Chat (Bai et al., 2023), LLaMA-2-7B-Instruct (Touvron et al., 2023), Qwen2-72B-Chat (Bai et al., 2023), LLaMA-2-7B-Instruct (Touvron et al., 2023), and LLaMA-3-70B-Instruct (Touvron et al., 2023; meta llama, 2024), we primarily use the hyperparameters from the official repositories. For the GPT model series, we access the models via API. All experiments are conducted on a machine equipped with 8 NVIDIA H100 GPUs, a 120-core CPU, and 960GB of memory.
Table 1. Comparison of PAS, BPO and not using APE (baseline). We can see PAS significantly outperform the baseline with an average improvement of 8 points. Additionally, when compared to the previous state-of-the-art model, BPO, our model still exceeds it by an average of 5.82 points.
Main Model APE-model Arena-hard Alpaca-Eval 2.0 Alpaca-Eval 2.0 (LC) Average
GPT-4-turbo-2024-04-09 None 76.6 46.12 55.02 59.25
GPT-4-1106-preview None 74.8 50 50 58.27
GPT-4-0613 None 37.9 15.8 30.2 27.97
GPT-3.5-turbo-1106 None 18.9 9.2 19.3 15.8
Qwen2-72B-Chat None 48.1 31.7 39.24 39.68
LLaMA-3-70b-instruct None 41.1 33.18 34.42 36.23
Average None 49.57 31.0 38.03 39.53
GPT-4-turbo-2024-04-09 BPO 76.60 54.65 55.28 62.18
GPT-4-1106-preview BPO 74.60 55.19 52.91 60.90
GPT-4-0613 BPO 38.60 19.61 34.08 30.76
GPT-3.5-turbo-1106 BPO 15.90 10.25 20.29 15.48
Qwen2-72b-Chat BPO 44.40 31.25 39.02 38.22
LLaMA-3-70b-instruct BPO 45.20 38.92 39.24 41.12
Average BPO 49.22 34.98 40.14 41.44
GPT-4-turbo-2024-04-09 PAS (PAS-None) 76.9 65.31 56.54 66.62 (+7.37)
GPT-4-1106-preview PAS (PAS-None) 78.8 65.92 53.63 66.12 (+7.85)
GPT-4-0613 PAS (PAS-None) 43.9 34.06 40.33 39.43 (+11.46)
GPT-3.5-turbo-1106 PAS (PAS-None) 22.1 15.82 23.31 20.41 (+4.61)
Qwen2-72B-Chat PAS (PAS-None) 52.2 45.53 44.31 47.35 (+7.67)
LLaMA-3-70b-instruct PAS (PAS-None) 50.3 45.01 40.52 45.28 (+9.05)
Average PAS (PAS-None) 54.03 45.37 43.20 47.53 (+8.00)
GPT-4-turbo-2024-04-09 PAS (PAS-BPO) 76.9 65.31 56.54 66.62 (+4.44)
GPT-4-1106-preview PAS (PAS-BPO) 78.8 65.92 53.63 66.12 (+5.22)
GPT-4-0613 PAS (PAS-BPO) 43.9 34.06 40.33 39.43 (+8.67)
GPT-3.5-turbo-1106 PAS (PAS-BPO) 22.1 15.82 23.31 20.41 (+4.93)
Qwen2-72B-Chat PAS (PAS-BPO) 52.2 45.53 44.31 47.35 (+9.13)
LLaMA-3-70b-instruct PAS (PAS-BPO) 50.3 45.01 40.52 45.28 (+4.16)
Average PAS (PAS-BPO) 54.03 45.37 43.20 47.53 (+6.09)
@@ 4.2. Main Experiments
To address Q1, we used Qwen2-7B-Chat as the base model due to its outstanding performance. We subsequently used the prompt complementary data to train a PAS model and compared it to both the baseline model without the APE model and the previous state-of-the-art (SoTA) APE model, BPO (Cheng et al., 2023). We integrated our model into multiple commonly used LLMs, including GPT-4-turbo-2024-04-09, GPT-4-1106-preview, GPT-4-0613, GPT-3.5-turbo-1106 (OpenAI, 2023a), Qwen2-72B-Chat (Bai et al., 2023), and LLaMA-3-70b-instruct (Touvron et al., 2023; meta llama, 2024).
@@
Moreover, when compared to the previous state-of-the-art model BPO, our model significantly outperforms it, resulting in an average improvement of 6.09 points. Each model achieves more than a 4-point average improvement across all six base models compared to BPO, with a notable increase of 9.13 points for Qwen2-72B-Chat, indicating a substantial improvement.
BPO is unstable and performs worse than the baseline in some cases, such as GPT-3.5-turbo-1106 and Qwen2-72B-Chat, indicating that the previous SoTA model cannot consistently outperform the baseline. Considering our model exceeds the baseline by 8.00 points and BPO by 6.09 points, it further demonstrates the effectiveness and robustness of our PAS model.
@@ 4.3. Effectiveness of PAS
Table 2. Comparison of PAS and BPO using the same base model, LLaMA-2-7b-instruct, shows that PAS significantly outperforms the BPO model across all LLMs.
Main Model Method Arena-hard Alpaca-Eval 2.0 Alpaca-Eval 2.0 (LC) Average
GPT-4-turbo-2024-04-09 BPO 76.60 54.65 55.28 62.18
GPT-4-1106-preview BPO 74.60 55.19 52.91 60.90
GPT-4-0613 BPO 38.60 19.61 34.08 30.76
GPT-3.5-turbo-1106 BPO 15.90 10.25 20.29 15.48
Qwen2-72b-Chat BPO 44.40 31.25 39.02 38.22
LLaMA-3-70b-instruct BPO 45.20 38.92 39.24 41.12
Average BPO 49.22 34.98 40.14 41.44
GPT-4-turbo-2024-04-09 PAS 73.54 62.58 54.03 63.38 (+1.20)
GPT-4-1106-preview PAS 75.52 64.06 53.07 64.22 (+3.32)
GPT-4-0613 PAS 40.13 33.11 36.70 36.65 (+5.89)
GPT-3.5-turbo-1106 PAS 18.02 16.18 23.67 19.29 (+3.81)
Qwen2-72B-Chat PAS 47.91 40.59 39.99 42.83 (+4.61)
LLaMA-3-70b-instruct PAS 46.30 43.17 38.77 42.74 (+1.62)
Average PAS 50.24 43.28 41.04 44.85 (+3.41)
@@ 4.4. Efficiency and Flexibility of PAS
To address Q3, we compare the data usage and flexibility of PAS with several SoTA models. This comparison highlights the efficiency and versatility of PAS.
@@ 4.6. Case Study
These case studies demonstrate that Automatic Prompt Engineering can significantly elevate the quality of AI interactions, making responses more contextually appropriate, logically consistent, and user-friendly.
@@ 4.7. Ablation Study
To address Q6, in this section, following Section 4.2, we first train a PAS model and conduct an ablation study by systematically replacing the data selection and data regeneration modules. Subsequently, we train another PAS model solely based on the data selection module. We compare the performance of these two models and summarize the results in Table 5.
From Table 5, it is evident that excluding the combined data selection and regeneration module leads to a significant decline in our model’s performance across all metrics. On average, our model’s performance decreased by 3.8 points, which is a notable reduction. Specifically, there was a decrease of 6.63 points in the GPT-4-1106-preview benchmark. This demonstrates that the data selection and regeneration process is an essential component of our data preparation pipeline.
Overall, the ablation study highlights the critical role of data selection and data regeneration in enhancing model performance. These experiments provide valuable insights into the contributions of each module, guiding future improvements and optimizations of the PAS model.
Table 5. Performance comparison between PAS models trained on a curated dataset and trained on a dataset without data selection.
Main Model PAS-model Arena-hard Alpaca-Eval 2.0 Alpaca-Eval 2.0 (LC) Average
GPT-4-turbo-2024-04-09 PAS 76.9 65.86 57.09 66.62
GPT-4-1106-preview PAS 78.8 65.92 53.63 66.12
GPT-4-0613 PAS 43.9 34.06 40.33 39.43
GPT-3.5-turbo-1106 PAS 22.1 15.82 23.31 20.41
Qwen2-72B-Chat PAS 52.2 45.53 44.31 47.35
LLaMA-3-70b-instruct PAS 50.3 45.01 40.52 45.28
Average PAS 54.03 45.37 43.20 47.53
GPT-4-turbo-2024-04-09 wo selection 75.0 57.97 49.52 60.83 (-5.79)
GPT-4-1106-preview wo selection 72.2 57.91 48.37 59.49 (-6.63)
GPT-4-0613 wo selection 38.7 31.59 36.19 35.49 (-3.94)
GPT-3.5-turbo-1106 wo selection 20.0 15.88 22.86 19.58 (-0.83)
Qwen2-72B-Chat wo selection 48.9 42.79 41.51 44.40 (-2.95)
LLaMA-3-70b-instruct wo selection 46.0 43.24 38.56 42.60 (-2.68)
Average wo selection 50.13 41.56 39.50 43.73 (-3.80)

### v2

@@ 2.3. Data Quality and Data Selection
Given the critical role of data quality and selection in enhancing model performance, our paper focuses on leveraging advanced data selection techniques to optimize prompts and their complementary data. By employing methods that integrate scoring mechanisms from LLMs and clustering techniques, we aim to efficiently identify and utilize high-quality data clusters for prompt engineering.
@@ 3.1. Prompts Data Collection
In this section, we introduce the prompt collection process. To generate high-quality (prompt, complementary prompt) pairs, we first need to select high-quality prompts. To achieve this, we use two curated datasets: the LMSYS-1M dataset (Zheng et al., 2023) and the WildChat dataset (Zhao et al., 2024). We use $P$ to denote the prompt dataset. As shown in Figure 3(a), our data selection process involves three main steps to ensure the quality and relevance of the data:
Deduplication
Deduplication is performed to create a diverse and efficient dataset. Using the SimCSE bge embedding model, all prompts from the LMSYS-1M and WildChat datasets are embedded.
$P_{\text{embed}}=\text{SimCSE}(P)$
Subsequently, the HNSW clustering algorithm is employed to group these embeddings. A small subset of data is then extracted from each cluster to reduce redundancy.
Quality Selection
Quality filtering is performed to select high-quality data because such data can not only reduce computational costs but also enhance the model’s performance. For quality selection, the BaiChuan 13b model (Yang et al., 2023b) serves as the base model. We filter out low-quality entries using the formula below:
$Q_{\text{score}}(p_{i})=\text{BaiChuan 13b}(p_{i})$
$P_{\text{filtered}}=\{p_{i}\in P\mid Q_{\text{score}}(p_{i})\geq\tau\}$
Here, $Q_{\text{score}}(p_{i})$ represents the quality score assigned by the BaiChuan 13b model to prompt $p_{i}$, and $\tau$ denotes the quality threshold. By employing quality selection, we aim to enhance the overall quality of prompt data.
Classification
To facilitate few-shot learning for prompt complementary dataset generation, we categorize the data into different categories. For accurate classification, we fine-tune a BaiChuan 13b model (Yang et al., 2023b) using 60,000 internally labeled classification data points from BaiChuan Inc. This process yields a classification model that categorizes prompts into common categories such as Q&A and coding.
The steps of deduplication, quality selection, and classification ensure diversity, quality, and accurate categorization of the data. Approximately 9,000 high-quality classified data points are obtained through these processes. Subsequently, these 9,000 high-quality prompts are utilized to generate high-quality (prompt, complementary prompt) pairs.
@@ 21 $isCorrect\leftarrow\text{FewShot Eval with }D_{\text{golden}}$;
To generate a high-quality prompt complementary dataset, we designed an automated data generation pipeline based on Few-Shot Learning. The algorithm mainly consists of two phases: data generation and data selection with regeneration. We use a set of golden data $D_{\text{golden}}=\{(p_{i},a_{i})\}_{i=1}^{N}$, containing 4 to 5 pairs of few-shot examples for each category from BaiChuan Inc. Then, we utilize the prompt dataset $P_{\text{golden}}$ from Section 3.1 to generate high-quality (prompt, complementary prompt) pairs.
@@
For each prompt $p_{j}\in P_{\text{golden}}$ in every category, we utilize the Few-Shot Learning method based on the prompt in Figure 4 to generate a corresponding complementary prompt $a_{j}$ based on the golden data $D_{\text{golden}}$. The generated prompt-complementary prompt pair $(p_{j},a_{j})$ is then added to $D_{\text{generated}}$.
@@ 3.3. Prompt Complementary Dataset
In this section, we present a detailed analysis of the generated prompt-complementary dataset. The dataset comprises approximately 9,000 high-quality (prompt, complementary prompt) pairs, as illustrated in Figure 6. Figure 6 summarizes the distribution of the dataset, revealing that it spans 14 categories, with each category containing approximately 500 data points. This wide coverage of various prompt cases underscores the dataset’s exceptional generalization capability. Moreover, there is a significant presence of Coding and Q&A data, which are widely utilized functions, thus justifying their prominent representation.
Despite these strengths in data distribution, our automated process for generating complementary prompts, as detailed in Section 3.2, allows us to exert control over the categories of generated data. This flexibility enables our method to cater to both general-purpose models and specialized data needs, thereby enhancing prompt adaptability across specific domains. Tailoring data generation to specific domains facilitates comprehensive training for diverse PAS tasks across various fields.
Overall, the broad coverage of various prompt cases demonstrates the dataset’s exceptional generalization ability, with emphasis on critical functionalities. Additionally, our method’s capability to control dataset distribution supports the development of PAS systems across all domains.
@@ 3.4. Automatic Prompt Complementary Plug-and-Play System PAS
Table 1. Comparison of PAS, BPO and not using APE (baseline). We can see PAS significantly outperform the baseline with an average improvement of 8 points. Additionally, when compared to the previous state-of-the-art model, BPO, our model still exceeds it by an average of 6.09 points.
Main Model APE-model Arena-hard Alpaca-Eval 2.0 Alpaca-Eval 2.0 (LC) Average $\uparrow$
GPT-4-turbo-2024-04-09 - 76.60 46.12 55.02 59.25 -
GPT-4-1106-preview - 74.80 50.00 50.00 58.27 -
GPT-4-0613 - 37.9 15.80 30.20 27.97 -
GPT-3.5-turbo-1106 - 18.90 9.20 19.30 15.80 -
Qwen2-72b-Instruct - 48.10 31.70 39.24 39.68 -
LLaMA-3-70b-Instruct - 41.10 33.18 34.42 36.23 -
Average - 49.57 31.0 38.03 39.53 -
GPT-4-turbo-2024-04-09 BPO 76.60 54.65 55.28 62.18 +2.93
GPT-4-1106-preview BPO 74.60 55.19 52.91 60.90 +2.63
GPT-4-0613 BPO 38.60 19.61 34.08 30.76 +2.79
GPT-3.5-turbo-1106 BPO 15.90 10.25 20.29 15.48 -0.32
Qwen2-72b-Instruct BPO 44.40 31.25 39.02 38.22 -1.46
LLaMA-3-70b-Instruct BPO 45.20 38.92 39.24 41.12 +1.59
Average BPO 49.22 34.98 40.14 41.44 +1.91
GPT-4-turbo-2024-04-09 PAS 76.90 65.31 56.54 66.62 +7.37
GPT-4-1106-preview PAS 78.80 65.92 53.63 66.12 +7.85
GPT-4-0613 PAS 43.90 34.06 40.33 39.43 +11.46
GPT-3.5-turbo-1106 PAS 22.10 15.82 23.31 20.41 +4.61
Qwen2-72b-Instruct PAS 52.20 45.53 44.31 47.35 +7.67
LLaMA-3-70b-Instruct PAS 50.30 45.01 40.52 45.28 +9.05
Average PAS 54.03 45.37 43.20 47.53 +8.00
@@ 4.1. Experiments Setting
For PAS models, we select several smaller models to efficiently train a PAS model. We select Qwen2-7b-Instruct (Bai et al., 2023), LLaMA-2-7b-Instruct (Touvron et al., 2023). Then we utilize our trained PAS models to some massive SoTA models, i.e. GPT-4-turbo-2024-04-09, GPT-4-1106-preview, GPT-4-0613, GPT-3.5-turbo-1106 (OpenAI, 2023a), Qwen2-72b-Instruct (Bai et al., 2023), and LLaMA-3-70b-Instruct (Touvron et al., 2023; meta llama, 2024).
@@
For Qwen2-7B-Instruct (Bai et al., 2023), LLaMA-2-7B-Instruct (Touvron et al., 2023), Qwen2-72B-Instruct (Bai et al., 2023), LLaMA-2-7B-Instruct (Touvron et al., 2023), and LLaMA-3-70B-Instruct (Touvron et al., 2023; meta llama, 2024), we primarily use the hyperparameters from the official repositories. For the GPT model series, we access the models via API. All experiments are conducted on a machine equipped with 8 NVIDIA H100 GPUs, a 120-core CPU, and 960GB of memory.
@@ 4.2. Main Experiments
To address Q1, we used Qwen2-7B-Instruct as the base model due to its outstanding performance. We subsequently used the prompt complementary data to train a PAS model and compared it to both the baseline model without the APE model and the previous state-of-the-art (SoTA) APE model, BPO (Cheng et al., 2023). We integrated our model into multiple commonly used LLMs, including GPT-4-turbo-2024-04-09, GPT-4-1106-preview, GPT-4-0613, GPT-3.5-turbo-1106 (OpenAI, 2023a), Qwen2-72b-Instruct (Bai et al., 2023), and LLaMA-3-70b-Instruct (Touvron et al., 2023; meta llama, 2024).
@@
Moreover, when compared to the previous state-of-the-art model BPO, our model significantly outperforms it, resulting in an average improvement of 6.09 points. Each model achieves more than a 4-point average improvement across all six base models compared to BPO, with a notable increase of 9.13 points for Qwen2-72b-Instruct, indicating a substantial improvement.
BPO is unstable and performs worse than the baseline in some cases, such as GPT-3.5-turbo-1106 and Qwen2-72b-Instruct, indicating that the previous SoTA model cannot consistently outperform the baseline. Considering our model exceeds the baseline by 8.00 points and BPO by 6.09 points, it further demonstrates the effectiveness and robustness of our PAS model.
@@
Table 2. Comparison of PAS and BPO using the same base model, LLaMA-2-7b-Instruct. The results demonstrate that PAS outperforms the BPO model consistently across all LLMs when using the same base model, LLaMA-2-7b-Instruct.
Main Model Method Arena-hard Alpaca-Eval 2.0 Alpaca-Eval 2.0 (LC) Average $\uparrow$
GPT-4-turbo-2024-04-09 BPO 76.60 54.65 55.28 62.18 -
GPT-4-1106-preview BPO 74.60 55.19 52.91 60.90 -
GPT-4-0613 BPO 38.60 19.61 34.08 30.76 -
GPT-3.5-turbo-1106 BPO 15.90 10.25 20.29 15.48 -
Qwen2-72b-Instruct BPO 44.40 31.25 39.02 38.22 -
LLaMA-3-70b-Instruct BPO 45.20 38.92 39.24 41.12 -
Average BPO 49.22 34.98 40.14 41.44 -
GPT-4-turbo-2024-04-09 PAS 73.54 62.58 54.03 63.38 +1.20
GPT-4-1106-preview PAS 75.52 64.06 53.07 64.22 +3.32
GPT-4-0613 PAS 40.13 33.11 36.70 36.65 +5.89
GPT-3.5-turbo-1106 PAS 18.02 16.18 23.67 19.29 +3.81
Qwen2-72b-Instruct PAS 47.91 40.59 39.99 42.83 +4.61
LLaMA-3-70b-Instruct PAS 46.30 43.17 38.77 42.74 +1.62
Average PAS 50.24 43.28 41.04 44.85 +3.41
@@ 4.4. Efficiency and Flexibility of PAS
To address Question 3 (Q3), we compare the data efficiency and flexibility of the PAS approach with SoTA APE models. This comparison underscores PAS’s effectiveness in terms of both data utilization and adaptability.
@@ 4.6. Case Study
These case studies demonstrate that our PAS system can significantly elevate the quality of AI interactions, making responses more contextually appropriate, logically consistent, and user-friendly.
Table 5. Performance comparison between PAS trained on a curated dataset and PAS trained without the Prompt Selection Module and Prompt Complementary Data Regeneration Module.
Main Model PAS-model Arena-hard Alpaca-Eval 2.0 Alpaca-Eval 2.0 (LC) Average $\uparrow$
GPT-4-turbo-2024-04-09 PAS 76.9 65.86 57.09 66.62 -
GPT-4-1106-preview PAS 78.8 65.92 53.63 66.12 -
GPT-4-0613 PAS 43.9 34.06 40.33 39.43 -
GPT-3.5-turbo-1106 PAS 22.1 15.82 23.31 20.41 -
Qwen2-72b-Instruct PAS 52.2 45.53 44.31 47.35 -
LLaMA-3-70b-Instruct PAS 50.3 45.01 40.52 45.28 -
Average PAS 54.03 45.37 43.20 47.53 -
GPT-4-turbo-2024-04-09 wo prompt selection 73.90 64.90 54.62 64.47 -2.15
GPT-4-1106-preview wo prompt selection 74.6 64.98 50.01 63.20 -2.92
GPT-4-0613 wo prompt selection 39.7 33.68 37.44 36.94 -2.49
GPT-3.5-turbo-1106 wo prompt selection 18.4 16.51 22.54 19.15 -1.26
Qwen2-72b-Instruct wo prompt selection 48.9 42.79 41.51 46.58 -0.77
LLaMA-3-70b-Instruct wo prompt selection 46.0 43.24 38.56 44.13 -1.15
Average wo prompt selection 50.97 45.20 41.07 45.75 -1.78
GPT-4-turbo-2024-04-09 wo regeneration 75.0 57.97 49.52 60.83 -5.79
GPT-4-1106-preview wo regeneration 72.2 57.91 48.37 59.49 -6.63
GPT-4-0613 wo regeneration 38.7 31.59 36.19 35.49 -3.94
GPT-3.5-turbo-1106 wo regeneration 20.0 15.88 22.86 19.58 -0.83
Qwen2-72b-Instruct wo regeneration 48.9 42.79 41.51 44.40 -2.95
LLaMA-3-70b-Instruct wo regeneration 46.0 43.24 38.56 42.60 -2.68
Average wo regeneration 50.13 41.56 39.50 43.73 -3.80
@@ 4.7. Ablation Study
To address Q6, in this section, following section 4.2, we first train a Qwen2-7b-Instruct to construct a PAS model using the curated dataset. Then we conduct two ablation studies. First, we replace the prompt data selection module with random prompt data selection and subsequently trained a PAS model without prompt selection (wo prompt selection). Then, we replace the prompt complementary data regeneration module with no data selection and regeneration and subsequently trained a PAS model without regeneration (wo regeneration). We compare the performance of these two models and summarize the results in Table 5.
Excluding Prompt Selection Module
From Table 5, it is evident that excluding the prompt data selection module (wo prompt selection) leads to a significant decline in our model’s performance across all metrics. On average, our model’s performance decreased by 1.78 points, which is a notable reduction. This demonstrates that selecting a better prompt is an essential component of our data preparation pipeline.
Excluding Prompt Complementary Data Regeneration Module
From Table 5, it is evident that excluding the combined data selection (wo regeneration) and regeneration module leads to a significant decline in our model’s performance across all metrics. On average, our model’s performance decreased by 3.8 points, which is a notable reduction. Specifically, there was a decrease of 6.63 points in the GPT-4-1106-preview benchmark. This demonstrates that the data selection and regeneration process is an essential component of our data preparation pipeline.
Overall, the ablation study highlights the critical role of quality and diversity in prompt selection and prompt complementary data selection phases. Both are critical in enhancing model performance. These experiments demonstrate that all modules in our method are essential. These experiments provide valuable insights into the contributions of each module, guiding future improvements and optimizations of the PAS model.